# Sim2Real in Reconstructive Spectroscopy: Deep Learning with Augmented Device-Informed Data Simulation

### v2

@@ III.2 Network Architecture: ReSpecNN
The ReSpecNN model is specifically tailored for our HDA scheme. In contrast to the previous model by Kim et al. Kim, Park, and Lee (2020), we incorporate an additional fully-connected block with skip connections to enhance the adaptability of our model in handling HDA-perturbed data. Instead of directly absorbing the information from the measured response matrix $\bm{R}$ into the calculations, which could be misleading due to inaccuracies in $\bm{R}$ measurements, we aim for the extra fully-connected block to potentially learn a more robust inverse of $\bm{R}$ through our HDA-perturbed data. This, in turn, enhances the overall model’s robustness, leading to improved reconstruction performance. Specifically, it’s important to note that the previous model by Kim et al. Kim, Park, and Lee (2020) initialize input data with information directly from measured $\bm{R}$ which might contain noise or misleading information. In comparison, through HDA, our ReSpecNN autonomously learns a more robust $\bm{R}$ for training.
@@
IV Results
In this section, we illustrate the performance of our Sim2Real approach on real-world data by comparing both the test accuracy and the inference time with a state-of-the-art optimization-based method: NNLS with Total Variation regularization (NNLS-TV)Sarwar et al. (2022).
Figure 6: Visualization of the reconstructed spectral signals using real-world data (Left: a single-peak sample, Right: a double-peak sample) produced by NNLS-TVSarwar et al. (2022) and our Sim2Real approach. The corresponding ground-truth spectra are measured by a commercial spectrometer. Our Sim2Real approach achieves a comparable performance with NNLS-TV. For any results with asymmetric peaks, the peak position is defined as the wavelength at which the maximum intensity occurs.
IV.1 Experimental Setup
@@
IV.2 Real-world Data Evaluation
To demonstrate the performance of our Sim2Real framework, we conducted evaluations using a real-world dataset, which is collected with a portable spectrometer deviceSarwar et al. (2022). This real-world dataset includes nine single-peak and five multiple-peak spectral signal samples, covering the wavelength range from $400$ nm to $650$ nm needed for the spectrometer’s targeted mobile application. The diverse peak profiles and the wide wavelength span ensure that the dataset is well-suited for testing the model under realistic conditions.
To evaluate the effectiveness of our proposed Sim2Real framework in comparison with the optimization-based method NNLS-TVSarwar et al. (2022), we focused mainly on two critical aspects: the accuracy of the spectral peak location and relative peak intensity. Instead of attempting to predict the entire spectral shape which would require a large number of deployed spectral encoders  [^1] In fact, the number of spectral encoders is the most important factor determining the spectral resolution, which in turn dictates the criterion of the Relative Peak Position Error. , our goal is to efficiently predict the most useful parameters for the mobile application, the spectral peak locations, using the least number of spectral encoders. Thus, we chose the following Relative Peak Position Error as our evaluation criterion, defined as
@@ IV.2 Real-world Data Evaluation
Figure 6 presents two spectrum reconstruction examples using our proposed deep neural network and the NNLS-TVSarwar et al. (2022) approach. When compared with the ground truths obtained by a commercial spectrometer, our approach exhibits a reliable performance on our miniature chip-scale spectrometer.
@@
For peak position predictions, Figure 7 show that our approach yields a result comparable with NNLS-TVSarwar et al. (2022) on aforementioned real-world datasetSarwar et al. (2022). In Figure 7, we calculated the Mean Absolute Error (MAE), defined as the average of the absolute values of the Relative Peak Position Error across all the spectral data samples. While offering a significantly faster inference speed (details to be discussed in the next section), our model can still maintain the relative errors within -5 to 5 percent.
@@
IV.3 Execution Time Analysis for Reconstruction Methods
@@
To demonstrate our superior inference time, we compared the execution time of our pre-trained model with the NNLS-TV solver using the real datasetSarwar et al. (2022). The results in Table 1 show that our model significantly reduces inference time. When conducting the execution time experiments in Table 1, we ensured fairness by utilizing only the CPU for both the NNLS-TV and our DL approach. Moreover, tests were performed on identical hardware setups to guarantee that the comparison solely reflects the computational efficiency of each method.
@@
Table 1: The average execution time (per sample, in milliseconds) for the NNLS-TVSarwar et al. (2022) method and our proposed approach on the real datasetSarwar et al. (2022). The NNLS-TV methodSarwar et al. (2022) was originally implemented with the MATLAB solver lsqnonneg in the paper, and to ensure a fair comparison with our model implemented in PyTorch, we also implemented their method using scipy.optimize.nnls in SciPy.
IV.4 Domain Shift in Reconstructive Spectroscopy
@@
As previously discussed in Section II.4, the domain shift usually exists between the device-informed simulated data $\bm{y}_{\textsf{sim}}$ and the real-world data $\bm{y}_{\textsf{real}}$. Many works only focus on improving the reconstruction accuracy within the datasets of the same distribution (i.e. training and testing exclusively on either simulated data or specific real-measured datasets) without attempting to bridge the domain gap between them. For instance, ResCNNKim, Park, and Lee (2020) has been demonstrated to achieve excellent spectral reconstruction performance and maintain stable results even under some noisy conditions. It preprocessed the input data from $\bm{y}$ to $\widetilde{\bm{x}}:=\bm{R}^{\dagger}\bm{y}$ ($\bm{R}^{\dagger}$ is the pseudoinverse of $\bm{R}$) for improved results. However, ResCNN trained only on simulated data does not perform optimally when directly applied to real-world dataSarwar et al. (2022).
@@
Figure 4 illustrates a comparison of the root mean square error (RMSE) under the Sim2Real setting between the spectral signals reconstructed by ResCNN and our ReSpecNN, against the ground truths for both the device-informed simulated data and the real data collected from our spectrometerSarwar et al. (2022). While ResCNN nearly perfectly fits the simulated data (with a lower RMSE), its performance on the real datasetSarwar et al. (2022) tends to drop significantly under the Sim2Real training setting, which confirms the domain shift phenomenon.
V Conclusion and Discussion
@@
In this paper, we introduce a novel Sim2Real framework for spectral signal reconstruction in spectroscopy, focusing on the sampling efficiency and inference time. Throughout the training process, only a single measurement of the response matrix $\bm{R}$ for the spectrometer device is required, with all other training data being simulated and generated. To address the domain shift between the real-world data and the device-informed simulated data, our Sim2Real framework introduces the hierarchical data augmentation approach to train our deep learning model. Furthermore, our neural network, ReSpecNN, which is trained exclusively on such augmented simulated data, is specifically designed for the reconstruction of real spectral signals. In our experiments, even with the simplest Gaussian noise augmentation, our Sim2Real method has achieved a 10-fold speed-up during the inference compared to the state-of-the-art optimization-based solver NNLS-TVSarwar et al. (2022), while demonstrating on par performance in terms of the solution quality. Although using only Gaussian noises for augmentation presents a limitation, the flexibility of our data augmentation suggests that improvements on the perturbation model of the response matrix $\bm{R}$ may further improve our Sim2Real framework by experimentally determining the noise distribution.
@@
Moving forward, to better handle outliers and further improve the robustness and accuracy of our model, we plan to apply the following approaches. First, it would be interesting to explore some improved noise patterns within our data augmentation process, for instance, incorporating the idea of adversarial training of deep networksGoodfellow, Shlens, and Szegedy (2015). Furthermore, while we focus on scenarios without labeled real training data, limited labeled data may be available in practice. These data can fine-tune our model, but the small size risks overfitting. To mitigate this while employing these data, we can use simulated training data for regularization during fine-tuning, as suggested in the recent study Ruiz et al. (2023).

### v3

@@ 3.2 Network Architecture: ReSpecNN
The ReSpecNN model is specifically tailored for our HDA scheme. In contrast to the previous model by Kim et al. [15], we incorporate an additional fully-connected block with skip connections to enhance the adaptability of our model in handling HDA-perturbed data. Instead of directly absorbing the information from the measured response matrix $\bm{R}$ into the calculations, which could be misleading due to inaccuracies in $\bm{R}$ measurements, we aim for the extra fully-connected block to potentially learn a more robust inverse of $\bm{R}$ through our HDA-perturbed data. This, in turn, enhances the overall model’s robustness, leading to improved reconstruction performance. Specifically, it’s important to note that the previous model by Kim et al. [15] initialize input data with information directly from measured $\bm{R}$ which might contain noise or misleading information. In comparison, through HDA, our ReSpecNN autonomously learns a more robust $\bm{R}$ for training.
@@
4 Results
In this section, we illustrate the performance of our Sim2Real approach on real-world data by comparing both the test accuracy and the inference time with a state-of-the-art optimization-based method: NNLS with Total Variation regularization (NNLS-TV)[13].
4.1 Experimental Setup
@@
Figure 6: Visualization of the reconstructed spectral signals using real-world data (Left: a single-peak sample, Right: a double-peak sample) produced by NNLS-TV[13] and our Sim2Real approach. The corresponding ground-truth spectra are measured by a commercial spectrometer. Our Sim2Real approach achieves a comparable performance with NNLS-TV. For any results with asymmetric peaks, the peak position is defined as the wavelength at which the maximum intensity occurs.
4.2 Real-world Data Evaluation
To demonstrate the performance of our Sim2Real framework, we conducted evaluations using a real-world dataset, which is collected with a portable spectrometer device[13]. This real-world dataset includes nine single-peak and five multiple-peak spectral signal samples, covering the wavelength range from $400$ nm to $650$ nm needed for the spectrometer’s targeted mobile application. The diverse peak profiles and the wide wavelength span ensure that the dataset is well-suited for testing the model under realistic conditions.
To evaluate the effectiveness of our proposed Sim2Real framework in comparison with the optimization-based method NNLS-TV[13], we focused mainly on two critical aspects: the accuracy of the spectral peak location and relative peak intensity. Instead of attempting to predict the entire spectral shape which would require a large number of deployed spectral encoders  [^1] In fact, the number of spectral encoders is the most important factor determining the spectral resolution, which in turn dictates the criterion of the Relative Peak Position Error. , our goal is to efficiently predict the most useful parameters for the mobile application, the spectral peak locations, using the least number of spectral encoders. Thus, we chose the following Relative Peak Position Error as our evaluation criterion, defined as
@@ 4.2 Real-world Data Evaluation
Figure 6 presents two spectrum reconstruction examples using our proposed deep neural network and the NNLS-TV[13] approach. When compared with the ground truths obtained by a commercial spectrometer, our approach exhibits a reliable performance on our miniature chip-scale spectrometer.
For peak position predictions, Figure 7 show that our approach yields a result comparable with NNLS-TV[13] on aforementioned real-world dataset[13]. In Figure 7, we calculated the Mean Absolute Error (MAE), defined as the average of the absolute values of the Relative Peak Position Error across all the spectral data samples. While offering a significantly faster inference speed (details to be discussed in the next section), our model can still maintain the relative errors within -5 to 5 percent.
@@
4.3 Execution Time Analysis for Reconstruction Methods
@@
To demonstrate our superior inference time, we compared the execution time of our pre-trained model with the NNLS-TV solver using the real dataset[13]. The results in Table 1 show that our model significantly reduces inference time. When conducting the execution time experiments in Table 1, we ensured fairness by utilizing only the CPU for both the NNLS-TV and our DL approach. Moreover, tests were performed on identical hardware setups to guarantee that the comparison solely reflects the computational efficiency of each method.
@@
Table 1: The average execution time (per sample, in milliseconds) for the NNLS-TV[13] method and our proposed approach on the real dataset[13]. The NNLS-TV method[13] was originally implemented with the MATLAB solver lsqnonneg in the paper, and to ensure a fair comparison with our model implemented in PyTorch, we also implemented their method using scipy.optimize.nnls in SciPy.
4.4 Domain Shift in Reconstructive Spectroscopy
@@
As previously discussed in Section 2.4, the domain shift usually exists between the device-informed simulated data $\bm{y}_{\textsf{sim}}$ and the real-world data $\bm{y}_{\textsf{real}}$. Many works only focus on improving the reconstruction accuracy within the datasets of the same distribution (i.e. training and testing exclusively on either simulated data or specific real-measured datasets) without attempting to bridge the domain gap between them. For instance, ResCNN[15] has been demonstrated to achieve excellent spectral reconstruction performance and maintain stable results even under some noisy conditions. It preprocessed the input data from $\bm{y}$ to $\widetilde{\bm{x}}:=\bm{R}^{\dagger}\bm{y}$ ($\bm{R}^{\dagger}$ is the pseudoinverse of $\bm{R}$) for improved results. However, ResCNN trained only on simulated data does not perform optimally when directly applied to real-world data[13].
@@
Figure 3 illustrates a comparison of the root mean square error (RMSE) under the Sim2Real setting between the spectral signals reconstructed by ResCNN and our ReSpecNN, against the ground truths for both the device-informed simulated data and the real data collected from our spectrometer[13]. While ResCNN nearly perfectly fits the simulated data (with a lower RMSE), its performance on the real dataset[13] tends to drop significantly under the Sim2Real training setting, which confirms the domain shift phenomenon.
5 Conclusion and Discussion
@@
In this paper, we introduce a novel Sim2Real framework for spectral signal reconstruction in spectroscopy, focusing on the sampling efficiency and inference time. Throughout the training process, only a single measurement of the response matrix $\bm{R}$ for the spectrometer device is required, with all other training data being simulated and generated. To address the domain shift between the real-world data and the device-informed simulated data, our Sim2Real framework introduces the hierarchical data augmentation approach to train our deep learning model. Furthermore, our neural network, ReSpecNN, which is trained exclusively on such augmented simulated data, is specifically designed for the reconstruction of real spectral signals. In our experiments, even with the simplest Gaussian noise augmentation, our Sim2Real method has achieved a 10-fold speed-up during the inference compared to the state-of-the-art optimization-based solver NNLS-TV[13], while demonstrating on par performance in terms of the solution quality. Although using only Gaussian noises for augmentation presents a limitation, the flexibility of our data augmentation suggests that improvements on the perturbation model of the response matrix $\bm{R}$ may further improve our Sim2Real framework by experimentally determining the noise distribution.
@@
Moving forward, to better handle outliers and further improve the robustness and accuracy of our model, we plan to apply the following approaches. First, it would be interesting to explore some improved noise patterns within our data augmentation process, for instance, incorporating the idea of adversarial training of deep networks[35]. Furthermore, while we focus on scenarios without labeled real training data, limited labeled data may be available in practice. These data can fine-tune our model, but the small size risks overfitting. To mitigate this while employing these data, we can use simulated training data for regularization during fine-tuning, as suggested in the recent study [36].